An electron is moving with a speed of 0.97c in a magnetic field of strength 10 to the power 8
G.
Calculate the peak frequency at which the electron will radiate.
1
Expert's answer
2020-04-20T10:41:44-0400

1 tesla = 10,000 gauss

here B=108 G which is equivalent to 104 tesla

frequency is "\frac{qB}{2\pi m} =\frac{1.6\times10^{-19}\times10^4}{2\times3.14\times9.1\times10^{-31}}=\frac{1.6\times10^{-15}}{0.565\times10^{-33}}=2.83\times10^{18}Hz"